# FIMP Dark Matter in bulk viscous non-standard cosmologies

### v2

@@ Abstract
non-standard cosmologies, dissipative cosmologies, FIMP dark matter candidates, particle physics and cosmology connection.
I Introduction
@@
Since the publication of the results obtained by the Supernova Search Team [1] and the Supernova Cosmology Project [2], which consistently provided evidence that the universe is currently undergoing an accelerating expansion phase, the $\Lambda$CDM model has become the most successful cosmological framework for describing the evolution of the universe and fitting observational cosmological data. This model aligns well with various cosmological observations, including Type Ia supernovae [3], measurements of the Hubble parameter [4, 5, 6], baryon acoustic oscillations [7], the cosmic microwave background [8, 9], among others. Regardless of its success, the model faces significant observational challenges, such as the $H_{0}$ tension, where local measurements of the Hubble constant $H_{0}$ using Cepheid host distance anchors (model-independent) show a discrepancy of $5\sigma$ with the value inferred from the Planck CMB assuming the $\Lambda$CDM model [10]. This tension is further supported by the H0LiCOW collaboration, which finds a discrepancy of $5.3\sigma$ with the value inferred from the Planck CMB [11]. From a more fundamental point of view, the nature and evolution of Dark Matter (DM) and Dark Energy (DE) remain open questions without definitive answers. The former, an as-yet-undetermined non-baryonic component of the universe, is estimated to be approximately five times more abundant than ordinary matter [9].
@@
All we know about DM is that, if it is a particle, it must be dark, neutral, non-baryonic matter, non-relativistic (Cold DM or CDM), and stable. Within this context, various theoretical frameworks propose potential DM candidates with these characteristics, such as Supersymmetry [12] and string theory [13]. Some of these DM particles are generally classified into two categories: Weakly Interacting Massive Particles (WIMPs) [14, 15, 16, 17, 18, 19], and Feebly Interacting Massive Particles (FIMPs) [20, 21, 22].
@@
WIMPs are massive particles, possibly in the GeV-TeV range. They are characterized by their interaction only through the electroweak force and gravity. The genesis mechanism for WIMPs in the early universe is the freeze-out process. This occurs when a non-negligible initial DM particle population is in thermal equilibrium with the Standard Model (SM) bath. As the universe expands and cools, the interactions between DM and SM particles gradually decrease, eventually ceasing and leaving behind the current observed DM relic density, with a value of $\Omega_{c}h^{2}=0.12$ according to Planck [9]. Understanding DM interactions is fundamental to identifying the most feasible methods for their detection. Among the most well-known experimental approaches to search for DM particles are direct detection [23], indirect detection [24], and collider experiments [25]. To be consistent with observations within the $\Lambda$CDM framework, the total thermally averaged DM production cross-section for this category must be approximately $\langle\sigma v\rangle_{0}=\text{few}\times 10^{-9}$ GeV [26]. Unfortunately, no favorable results have been obtained to date in the search for WIMP DM, which has led to the exploration of alternative production mechanisms for these elusive particles. In this context, one of the most recently studied mechanisms is freeze-in, which gives rise to the second category of DM candidates, the FIMPs.
@@
FIMPs are massive particles in the keV-TeV range. These particles interact so weakly with SM particles that they never reach thermal equilibrium with the surrounding medium in the early universe. Consequently, these particles undergo the Freeze-In process to establish their population. In this mechanism, the DM abundance is gradually built up as the universe cools, until the production ceases, leaving behind the current observed DM relic density. Even though the feeble interactions of FIMPs make it challenging to detect them with current experimental instruments, they are of special interest because these candidates, which are not being generated through thermal equilibrium, allow us to explore a less restricted parameter space compared to WIMPs. Following this purpose, many studies suggest that these candidates could be found in colliders and direct detection experiments as they can couple to long-lived particles due to the feeble couplings or may also be produced from the decay of heavy particles [27, 28, 29]. Furthermore, due to the mass range of FIMPs, they can be the origin of Warm DM (WDM) [30, 31, 32, 22, 33, 34, 35]. The possibility of this type of DM has an impact on the structure formation in the universe. While the CDM can explain the observed structure on scales above $\sim 1Mpc$, it faces challenges in explaining small-scale structure observations [36, 37], such as the missing satellite problem [38, 39]. This issue can be alleviated by WDM instead of a CDM [40], a possibility that is supported by many studies related to the number of satellites in the Milky Way and small halos with dwarf galaxies [37, 41, 42]. Therefore, FIMPs are more versatile than WIMPs, as they are less constrained and can alleviate other cosmological tensions.
@@
The DM relic density depends on the era in which its abundance is set. For instance, in the $\Lambda$CDM model, it is assumed that this relic density is established during the radiation-dominated era. Some interesting cosmological scenarios include an additional field ($\phi$) in the early universe, which generates different domination eras due to the decay of the field into SM particles, leading to entropy injection into the SM bath. This behavior expands the parameter space available for detecting DM particles and can modify DM production if it occurs during these new domination eras. These alternative scenarios, known as Non-Standard Cosmologies (NSCs), offer new avenues for DM detection and may re-open parameter spaces previously excluded in the $\Lambda$CDM model, potentially revealing viable DM candidates and mechanisms within these non-standard frameworks. Therefore, if DM is experimentally detected, its particle physics properties and characteristics—such as mass, interactions with SM particles, and coupling constants—must be determined in a way that is consistent with the $\Lambda$CDM framework. Conversely, it is necessary to explore new NSC scenarios where these quantities align with the cosmological model.
@@
In the effort to address tensions within the $\Lambda$CDM model, one approach is to consider that this model is a particular approximation of a more general cosmological framework. In this context, a natural extension is to incorporate non-perfect fluids, which offer a more realistic description of the cosmic medium [43]. This is especially relevant given that, in the $\Lambda$CDM model, all matter components of the universe are typically modeled as perfect fluids. When non-perfect fluids are taken into account, phenomena like viscosity arise, whose effects play a crucial role in many cosmological processes, including the reheating of the universe, the decoupling of neutrinos from the cosmic plasma, nucleosynthesis, and others. Even more, the effects of viscosity are significant in various astrophysical mechanisms, such as the collapse of radiating stars into neutron stars or black holes, and the accretion of matter around compact objects [43]. More recently and following the main idea, viscosity has been proposed as an alternative to mitigate some of the recent tensions in the $\Lambda$CDM model. For instance, a decaying DM scenario increases the expansion rate relative to $\Lambda$CDM, potentially alleviating the $H_{0}$ tension [44], with other viable solutions for this tension considering viscosity studied in Refs. [45, 46, 47] (for a comprehensive review of viscous cosmology in the early and late universe, see Ref. [48]).
@@
Working with non-perfect fluids requires a relativistic thermodynamic theory out of equilibrium, such as the Eckart [49, 50] or Israel-Stewart [51, 52] formalisms. Although Eckart’s theory is non-causal [53], it remains widely studied due to its mathematical simplicity compared to the full Israel-Stewart formalism. Nevertheless, Eckart’s approach serves as a convenient starting point for understanding dissipative effects in the universe, as the Israel-Stewart theory reduces to Eckart’s theory when the relaxation time for transient viscous effects is negligible [54]. On the other hand, viscosity can manifest in two forms: shear viscosity and bulk viscosity. While the former can play a significant role in certain contexts [55], our focus will be on bulk viscosity, which can emerge due to the presence of mixtures in the universe [56], the decaying of DM into relativistic particles [57, 58], the energy transfer between Cold Dark Matter (CDM) and the radiation fluid [59], and might exist in a hidden sector, reproducing various observational properties of disk galaxies [60, 61]. The recent advancements in gravitational wave detection have further opened up the possibility of observing dissipative effects in DM and DE through the dispersion and dissipation experienced by these waves as they propagate through a non-perfect fluid [62]. Additionally, bulk viscosity could contribute significantly to the emission of gravitational waves during neutron star mergers [63].
@@
Bulk viscous effects in the matter components of the universe have been deeply and widely studied in the literature, particularly about viscous DE components [64, 65, 66, 67, 68, 69]. However, a more common approach is to study a DM component that undergoes dissipative processes during its cosmic evolution [70, 71, 72, 73, 74, 75, 76, 77]. This approach is of particular interest because it can account for the recent accelerated expansion of the universe without requiring an explicit DE component, leading to unified DM models [78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88]. Previous studies have explored dissipative stiff matter fluids within the framework of the full Israel-Stewart theory [89]. Furthermore, bulk viscous DM has been analyze in various contexts, including inflation [90, 91, 54, 92], interacting fluids [56, 93, 94], and modified gravity [95, 68]. These effects have also been studied concerning singularities such as the Big Rip and Little Rip, across both classical and quantum regimes [64, 66, 96, 68, 97, 98, 99, 100, 101, 102, 103]. Further research has investigated the role of bulk viscosity in the radial oscillations of relativistic stars as well as its cosmological implications for Quark-Gluon plasma-filled universes [104, 105].
@@
In Ref. [106] the authors propose a novel NSC scenario in which the early universe is dominated by two interacting fluids, namely the new field $\phi$ and radiation, by considering that $\phi$ experiences dissipative processes during their cosmic evolution in the form of a bulk viscosity. Working in the framework of Eckart’s theory for non-perfect fluids, the authors consider an interaction term of the form $\Gamma_{\phi}\rho_{\phi}$, where $\Gamma_{\phi}$ represents the decay rate of the field and $\rho_{\phi}$ denotes its energy density, and a bulk viscosity described by $\xi=\xi_{0}\rho_{\phi}^{1/2}$. The latter has the advantage that, when the field fully decays, the dissipation is negligible and the standard $\Lambda$CDM cosmology is fully recovered. This novel NSC scenario was applied to the study of the parameter space for WIMP DM candidate production, showing that it is possible to explore even wider regions in which the model can account for the actual DM relic abundance in the universe. This result is obtained by comparing the classical NSC scenario with their bulk viscous counterpart, considering the same values for the NSC parameters, namely, $\kappa$ (ratio between the initial energy density of $\phi$ and radiation), $T_{\text{end}}$ (the temperature at which $\phi$ decays), $\xi_{0}$, and $\omega$ (the barotropic index of $\phi$); and for the WIMP DM candidate $m_{\chi}$ (the DM mass) and $\langle\sigma v\rangle$. From the study, it can be noted that in the bulk viscous NSC, there is a boost in the production of the field $\phi$ in comparison with the NSC case and, therefore, an increased injection of entropy to the SM bath. On the other hand, for certain values of the parameter space $(T_{\text{end}},\kappa)$, it can be observed that the model with viscosity allows a considerable range of lower values of $\kappa$ to reproduce the DM relic density compared with the NSC. The latter means that for larger values of $\kappa$ both cases, NSC with and without bulk viscosity, are comparable due to the contribution of the viscosity being almost neglected, and the addition of viscosity enables a bit larger values of $T_{\text{end}}$. Also, in the $(m_{\chi},\langle\sigma v\rangle)$ space for the WIMP DM candidate in the same range of masses, NSC and NSC with bulk viscosity allow higher values of $m_{\chi}$ unlike the $\Lambda$CDM model, and the effects of dissipative processes let us to explore smaller total thermally averaged DM production cross-section for the candidate. These results raise the following question: How do dissipative processes influence the parameter space for FIMP DM production?
@@
The aim of this paper is to revisit the bulk viscous NSC studied in [106] by investigating the impact of these dissipative effects on the production and relic density of FIMP DM candidates. Following the same scheme as in [106] for WIMP DM candidates, we investigate the parameter space that can accurately reproduce the current DM relic density by adjusting both the model and DM parameters within this novel NSC framework for FIMP DM candidates.
@@
This paper is organized as follows: Section II provides a brief review of the original NSC scenario. Subsection II.1 examines its applicability to FIMP DM candidates. In section III, we introduce the bulk viscous extension to the original NSC model, being compared the two models for FIMP DM candidates with a constant total thermally averaged DM production cross-section in subsection III.1. The analysis of the parameter space for DM production that results in the current relic density is presented in Subsection III.2. In Subsection III.3, we briefly study this bulk viscous NSC for FIMP DM candidates with a non-constant total thermally averaged DM production cross-section. Finally, in Section IV, we present some conclusions and future directions. Throughout this paper, we use $c=1$ units.
II Non-standard cosmologies

### v3

# non-standard cosmologies; dissipative cosmologies; FIMP dark matter candidates; particle physics and cosmology connection

@@ Abstract
non-standard cosmologies; dissipative cosmologies; FIMP dark matter candidates; particle physics and cosmology connection
[…]
1 Introduction
@@
Since the publication of the results obtained by the Supernova Search Team Riess et al. (1998) and the Supernova Cosmology Project Perlmutter et al. (1999), which consistently provided evidence that the universe is currently undergoing an accelerating expansion phase, the $\Lambda$CDM model has become the most successful cosmological framework for describing the evolution of the universe and fitting observational cosmological data. This model aligns well with various cosmological observations, including Type Ia supernovae Scolnic et al. (2022), measurements of the Hubble parameter Moresco et al. (2012); Zhang et al. (2014); Moresco (2015), baryon acoustic oscillations Eisenstein et al. (2005), the cosmic microwave background Bennett et al. (2013); Aghanim et al. (2020), among others. Regardless of its success, the model faces significant observational challenges, such as the $H_{0}$ tension, where local measurements of the Hubble constant $H_{0}$ using Cepheid host distance anchors (model-independent) show a discrepancy of $5\sigma$ with the value inferred from the Planck CMB assuming the $\Lambda$CDM model Riess et al. (2022). This tension is further supported by the H0LiCOW collaboration, which finds a discrepancy of $5.3\sigma$ with the value inferred from the Planck CMB Wong et al. (2020). From a more fundamental point of view, the nature and evolution of Dark Matter (DM) and Dark Energy (DE) remain open questions without definitive answers. The former, an as-yet-undetermined non-baryonic component of the universe, is estimated to be approximately five times more abundant than ordinary matter Aghanim et al. (2020).
@@
All we know about DM is that, if it is a particle, it must be dark, neutral, non-baryonic matter, non-relativistic (Cold DM or CDM), and stable. Within this context, various theoretical frameworks propose potential DM candidates with these characteristics, such as Supersymmetry Jungman et al. (1996) and string theory King and Roberts (2007). Some of these DM particles are generally classified into two categories: Weakly Interacting Massive Particles (WIMPs) Steigman and Turner (1985); Bertone et al. (2005); Arcadi et al. (2018); Roszkowski et al. (2018); Arcadi et al. (2024); Singh et al. (2024), and Feebly Interacting Massive Particles (FIMPs) Bernal et al. (2017); Chu et al. (2012); Hall et al. (2010).
@@
WIMPs are massive particles, possibly in the GeV-TeV range. They are characterized by their interaction only through the electroweak force and gravity. The genesis mechanism for WIMPs in the early universe is the freeze-out process. This occurs when a non-negligible initial DM particle population is in thermal equilibrium with the Standard Model (SM) bath. As the universe expands and cools, the interactions between DM and SM particles gradually decrease, eventually ceasing and leaving behind the current observed DM relic density, with a value of $\Omega_{c}h^{2}=0.12$ according to Planck Aghanim et al. (2020). Understanding DM interactions is fundamental to identifying the most feasible methods for their detection. Among the most well-known experimental approaches to search for DM particles are direct detection Abi et al. (2020), indirect detection Feng (2010), and collider experiments Alves et al. (2008). To be consistent with observations within the $\Lambda$CDM framework, the total thermally averaged DM production cross-section for this category must be approximately $\langle\sigma v\rangle_{0}=\text{few}\times 10^{-9}$ GeV Steigman et al. (2012). Unfortunately, no favorable results have been obtained to date in the search for WIMP DM, which has led to the exploration of alternative production mechanisms for these elusive particles. In this context, one of the most recently studied mechanisms is freeze-in, which gives rise to the second category of DM candidates, the FIMPs.
@@
FIMPs are massive particles in the keV-TeV range. These particles interact so weakly with SM particles that they never reach thermal equilibrium with the surrounding medium in the early universe. Consequently, these particles undergo the Freeze-In process to establish their population. In this mechanism, the DM abundance is gradually built up as the universe cools, until the production ceases, leaving behind the current observed DM relic density. These DM candidates can emerge from various particle physics models that require the introduction of new symmetries beyond the SM. One example is the Higgs portal mechanism, which incorporates a scalar field as the DM component, stabilized by a $Z_{2}$ odd symmetry to prevent its decay. Alternatively, in scenarios where the DM candidate is fermionic, the associated vector boson corresponds to a gauge boson of a hidden gauge symmetry, such as $SU(2)$ Bernal et al. (2017). Another well-motivated DM candidate is the pseudo-Nambu-Goldstone boson, exemplified by Majoron DM, which arises when the $U(1)_{B-L}$ symmetry is spontaneously broken Frigerio et al. (2011). Additional scenarios include the spontaneous breaking of $Z_{4}$ symmetry within an inverse seesaw model, leading to neutrino mass generation alongside a FIMP DM candidate Wang et al. (2024), or the implementation of the Froggatt-Nielsen mechanism via the introduction of a $U(1)_{FN}$ symmetry Babu et al. (2023). Even though the feeble interactions of FIMPs make it challenging to detect them with current experimental instruments, they are of special interest because these candidates, which are not being generated through thermal equilibrium, allow us to explore a less restricted parameter space compared to WIMPs. Following this purpose, many studies suggest that these candidates could be found in colliders and direct detection experiments as they can couple to long-lived particles due to the feeble couplings or may also be produced from the decay of heavy particles Junius (2022); Barman et al. (2024); Bélanger et al. (2024); Blinnikov and Khlopov (1982, 1983); Khlopov et al. (1991). Furthermore, due to the mass range of FIMPs, they can be the origin of Warm DM (WDM) D’Eramo and Lenoci (2021); McDonald (2016); Klasen and Yaguna (2013); Hall et al. (2010); Frigerio et al. (2011); de Giorgi and Vogl (2023); Kuo et al. (2018). The possibility of this type of DM has an impact on the structure formation in the universe. While the CDM can explain the observed structure on scales above $\sim 1Mpc$, it faces challenges in explaining small-scale structure observations Bode et al. (2001); de Vega and Sanchez (2011), such as the missing satellite problem Klypin et al. (1999); Moore et al. (1999). This issue can be alleviated by WDM instead of a CDM Weinberg et al. (2015), a possibility that is supported by many studies related to the number of satellites in the Milky Way and small halos with dwarf galaxies de Vega and Sanchez (2011); Viel et al. (2005); Newton et al. (2021). Therefore, FIMPs are more versatile than WIMPs, as they are less constrained and can alleviate other cosmological tensions.
@@
The DM relic density depends on the era in which its abundance is set. For instance, in the $\Lambda$CDM model, it is assumed that this relic density is established during the radiation-dominated era. Some interesting cosmological scenarios include an additional field ($\phi$) in the early universe (scalar or fermion), which generates different domination eras due to the decay of the field into SM particles, leading to entropy injection into the SM bath. This behavior expands the parameter space available for detecting DM particles and can modify DM production if it occurs during these new domination eras. These alternative scenarios, known as Non-Standard Cosmologies (NSCs), offer new avenues for DM detection and may re-open parameter spaces previously excluded in the $\Lambda$CDM model, potentially revealing viable DM candidates and mechanisms within these non-standard frameworks. Therefore, if DM is experimentally detected, its particle physics properties and characteristics—such as mass, interactions with SM particles, and coupling constants—must be determined in a way that is consistent with the $\Lambda$CDM framework. Conversely, it is necessary to explore new NSC scenarios where these quantities align with the cosmological model.
@@
In the effort to address tensions within the $\Lambda$CDM model, one approach is to consider that this model is a particular approximation of a more general cosmological framework. In this context, a natural extension is to incorporate non-perfect fluids, which offer a more realistic description of the cosmic medium Maartens (1996). This is especially relevant given that, in the $\Lambda$CDM model, all matter components of the universe are typically modeled as perfect fluids. When non-perfect fluids are taken into account, phenomena like viscosity arise, whose effects play a crucial role in many cosmological processes, including the reheating of the universe, the decoupling of neutrinos from the cosmic plasma, nucleosynthesis, and others. Even more, the effects of viscosity are significant in various astrophysical mechanisms, such as the collapse of radiating stars into neutron stars or black holes, and the accretion of matter around compact objects Maartens (1996). More recently and following the main idea, viscosity has been proposed as an alternative to mitigate some of the recent tensions in the $\Lambda$CDM model. For instance, a decaying DM scenario increases the expansion rate relative to $\Lambda$CDM, potentially alleviating the $H_{0}$ tension Pandey et al. (2020), with other viable solutions for this tension considering viscosity studied in Refs. Yang et al. (2019); Di Valentino et al. (2021); Normann and Brevik (2021) (for a comprehensive review of viscous cosmology in the early and late universe, see Ref. Brevik et al. (2017)). It is important to highlight that these tensions (particularly the late-time tensions) are mentioned solely as a motivation to explore alternatives to the $\Lambda$CDM model by considering bulk viscous fluids. In this sense, we also include this extension in the early universe to study its impact on FIMP dark matter production as a potential dark matter candidate. Nonetheless, it is important to note that our model is not incompatible with these late-time approaches. Therefore, one could develop a comprehensive model that could alleviate the issues of the $\Lambda$CDM model at both early and late times.
@@
Working with non-perfect fluids requires a relativistic thermodynamic theory out of equilibrium, such as the Eckart Eckart (1940a, b) or Israel-Stewart Stewart (1977); Israel and Stewart (1979) formalisms. Although Eckart’s theory is non-causal Israel (1976), it remains widely studied due to its mathematical simplicity compared to the full Israel-Stewart formalism. Nevertheless, Eckart’s approach serves as a convenient starting point for understanding dissipative effects in the universe, as the Israel-Stewart theory reduces to Eckart’s theory when the relaxation time for transient viscous effects is negligible Maartens (1995). On the other hand, viscosity can manifest in two forms: shear viscosity and bulk viscosity. While the former can play a significant role in certain contexts Floerchinger et al. (2015), our focus will be on bulk viscosity, which can emerge due to the presence of mixtures in the universe Zimdahl (1996), the decaying of DM into relativistic particles Wilson et al. (2007); Mathews et al. (2008), the energy transfer between CDM and the radiation fluid Hofmann et al. (2001), and might exist in a hidden sector, reproducing various observational properties of disk galaxies Foot and Vagnozzi (2015, 2016). The recent advancements in gravitational wave detection have further opened up the possibility of observing dissipative effects in DM and DE through the dispersion and dissipation experienced by these waves as they propagate through a non-perfect fluid Goswami et al. (2017). Additionally, bulk viscosity could contribute significantly to the emission of gravitational waves during neutron star mergers Alford et al. (2018). It is interesting to remark that until now there has been no fundamental origin for bulk viscosity from particle physics; it is instead treated as an effective description.
@@
Bulk viscous effects in the matter components of the universe have been deeply and widely studied in the literature, particularly about viscous DE components Nojiri and Odintsov (2004); Capozziello et al. (2006); Nojiri and Odintsov (2005); Cataldo et al. (2005); Brevik (2006); Cruz et al. (2017). However, a more common approach is to study a DM component that undergoes dissipative processes during its cosmic evolution Velten and Schwarz (2012); Acquaviva and Beesham (2014); Cruz et al. (2017, 2018, 2022, 2023); Gómez et al. (2023); Cruz et al. (2023). This approach is of particular interest because it can account for the recent accelerated expansion of the universe without requiring an explicit DE component, leading to unified DM models Fabris et al. (2006); Avelino and Nucamendi (2009); Li and Barrow (2009); Avelino and Nucamendi (2010); Hipolito-Ricaldi et al. (2010); Velten and Schwarz (2011); Gagnon and Lesgourgues (2011); Bruni et al. (2013); Cruz et al. (2017, 2020, 2021). Previous studies have explored dissipative stiff matter fluids within the framework of the full Israel-Stewart theory Mak and Harko (2004). Furthermore, bulk viscous DM has been analyze in various contexts, including inflation Padmanabhan and Chitre (1987); Barrow (1988); Maartens (1995); Maartens and Mendez (1997), interacting fluids Zimdahl (1996); Avelino et al. (2013); Hernández-Almada et al. (2020), and modified gravity Brevik et al. (2005); Brevik (2006). These effects have also been studied concerning singularities such as the Big Rip and Little Rip, across both classical and quantum regimes Nojiri and Odintsov (2004, 2005); Brevik and Gorbunova (2005); Brevik (2006); Brevik and Gorbunova (2008); Brevik et al. (2010a, b, 2011); Contreras et al. (2016, 2018); Cruz et al. (2022). Further research has investigated the role of bulk viscosity in the radial oscillations of relativistic stars as well as its cosmological implications for Quark-Gluon plasma-filled universes Barta (2019); Bravo Medina et al. (2019).
@@
In Ref. González et al. (2024) the authors propose a novel NSC scenario in which the early universe is dominated by two interacting fluids, namely the new field $\phi$ and radiation, by considering that $\phi$ experiences dissipative processes during their cosmic evolution in the form of a bulk viscosity. Working in the framework of Eckart’s theory for non-perfect fluids, the authors consider an interaction term of the form $\Gamma_{\phi}\rho_{\phi}$, where $\Gamma_{\phi}$ represents the decay rate of the field and $\rho_{\phi}$ denotes its energy density, and a bulk viscosity described by $\xi=\xi_{0}\rho_{\phi}^{1/2}$. The latter has the advantage that, when the field fully decays, the dissipation is negligible and the standard $\Lambda$CDM cosmology is fully recovered. This novel NSC scenario was applied to the study of the parameter space for WIMP DM candidate production, showing that it is possible to explore even wider regions in which the model can account for the actual DM relic abundance in the universe. This result is obtained by comparing the classical NSC scenario with their bulk viscous counterpart, considering the same values for the NSC parameters, namely, $\kappa$ (ratio between the initial energy density of $\phi$ and radiation), $T_{\text{end}}$ (the temperature at which $\phi$ decays), $\xi_{0}$, and $\omega$ (the barotropic index of $\phi$); and for the WIMP DM candidate $m_{\chi}$ (the DM mass) and $\langle\sigma v\rangle$. From the study, it can be noted that in the bulk viscous NSC, there is a boost in the production of the field $\phi$ in comparison with the NSC case and, therefore, an increased injection of entropy to the SM bath. On the other hand, for certain values of the parameter space $(T_{\text{end}},\kappa)$, it can be observed that the model with viscosity allows a considerable range of lower values of $\kappa$ to reproduce the DM relic density compared with the NSC. The latter means that for larger values of $\kappa$ both cases, NSC with and without bulk viscosity, are comparable due to the contribution of the viscosity being almost neglected, and the addition of viscosity enables a bit larger values of $T_{\text{end}}$. Also, in the $(m_{\chi},\langle\sigma v\rangle)$ space for the WIMP DM candidate in the same range of masses, NSC and NSC with bulk viscosity allow higher values of $m_{\chi}$ unlike the $\Lambda$CDM model, and the effects of dissipative processes let us to explore smaller total thermally averaged DM production cross-section for the candidate. These results raise the following question: How do dissipative processes influence the parameter space for FIMP DM production?
@@
The aim of this paper is to revisit the bulk viscous NSC studied in González et al. (2024) by investigating the impact of these dissipative effects on the production and relic density of FIMP DM candidates. Following the same scheme as in González et al. (2024) for WIMP DM candidates, we investigate the parameter space that can accurately reproduce the current DM relic density by adjusting both the model and DM parameters within this novel NSC framework for FIMP DM candidates.
@@
This paper is organized as follows: Section 2 provides a brief review of the original NSC scenario. Subsection 2.1 examines its applicability to FIMP DM candidates. In section 3, we introduce the bulk viscous extension to the original NSC model, being compared the two models for FIMP DM candidates with a constant total thermally averaged DM production cross-section in subsection 3.1. The analysis of the parameter space for DM production that results in the current relic density is presented in Subsection 3.2. In Subsection 3.3, we briefly study this bulk viscous NSC for FIMP DM candidates with a non-constant total thermally averaged DM production cross-section. Finally, in Section 4, we present some conclusions and future directions. Throughout this paper, we use $c=1$ units.
2 Non-standard cosmologies